OpenStudy (anonymous):

what is a equation to a line that is paraell to (12,6) and (37,13) and passes through (8,12)

OpenStudy (bahrom7893):

parallel => slope is the same.

OpenStudy (anonymous):

So find the slope\[m=(y _{2}-y _{1})/(x _{2}-x _{1})=(13-6)/(37-12)=7/25\]

OpenStudy (anonymous):

Now that we have the slope, we can use it and the required point in the point-slope form to get an equation of the line. Recall \[y−y_1=m(x−x_1)\] so, \[y−12=(7/25)(x−8)\] Since the form of the answer wasn't specified, I'd just leave it like that.